OpenStudy (anonymous):

if 2(a/k)+1=9 what does a/k equal?

OpenStudy (ddcamp):

Just say that a/k equals some other variable x, and substitute: 2x + 1 = 9 2x + 1 - 1 = 9-1 2x = 8 2x ÷ 2 = 8÷2 x = 4 a/k = 4

OpenStudy (anonymous):